OpenStudy (anonymous):

Find the perimeter and area of a right triangle if one leg measures 8 cm and the other leg measures 15 cm. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.

OpenStudy (anonymous):

hypen=\[15^{2}+8^{2}=\sqrt{289}\] =17 P = 8+15+17=50 A= (b*h)/2 =(15*8)/2 =60

OpenStudy (anonymous):

If I deserve medal then give me one